(Decide whether the ordered pair is a solution of the system of linear equations) Help me ! How to do ?
(1;1) (-1;0)
2x+y=3
x-2y=-1

Answer by askmemath(368)

Simply substitute the points in place of X and Y in the equations. If the points satisfy the equation, then it is said to be a solution
For e.g. (1,1) and 2x+y = 3
Substituting
Left Hand side
2(1)+ 1
= 2+1
= 3
= Right Hand side
Thus (1,1) is a solution of the equation 2x+y = 3
Can you solve for the rest?